OpenStudy (anonymous):

A bouncing ball reaches a height of 27 feet at its first peak, 18 feet at its second peak, and 12 feet at its third peak. Describe how a sequence can be used to determine the height of the ball when it reaches its fourth peak. Pleaaaase I need help!!

OpenStudy (dan815):

27*2/3=18 18*2/3=12

OpenStudy (phi):

I would look at the ratio of two terms and the difference of two terms to see if there might be a pattern for example, 27, 18, 12 the difference between 27 and 18 is 9 the difference between 18 and 12 is 6 That is not a constant difference. Let's try ratios: 27/18 = 1.5 18/12 = 1.5 that looks promising... we get the same ratio so on the next bounce, expect 18/x = 1.5 maybe it's easier if we flip the ratio 1.5 (3/2) to 2/3 and so 18= 2/3 * 27 12= 2/3* 18 = 2/3 * 2/3 * 27 If we try to make a pattern first peak = (2/3)^0 * 27 2nd peak = (2/3)^1 *27 3rd peak = (2/3)^2 * 27

OpenStudy (anonymous):

Thank you!!!